The circular surface of a table has a diameter of 4 ft. What will cost to have the top refinished if the company charges $5 per square foot for the refinishing? Use 3.14 for and round your answer to the nearest cent. Show all the work and write the answer in a complete sentence.
Answer by checkley77(12844)

AREA=PI*R^2
AREA=3.14*(4/2)^2
AREA=3.14*2^2
AREA=3.14*4
AREA=12.56 FT^2. ANS.
12.56*5=$62.50 IS THE COST TO REFINISH THE TABLE TOP.